# Non-Unitary Quantum Many-Body Dynamics using the Faber Polynomial Method

### v2

@@ 1 Introduction
In recent years, the scientific community has shown a growing interest in elucidating the distinctive characteristics of many-body quantum systems subjected to effective non-unitary dynamics. In quantum mechanics, non-unitary dynamics typically arises when a closed quantum system interacts with an external environment, leading to dissipation, decoherence, or wave function collapse that disrupts the usual unitary Schrödinger evolution. While a fully microscopic description of both environment and system is a daunting task, in many cases of experimental and theoretical interest one can assume the dynamics of the environment to be sufficiently fast, allowing for the derivation of a local-in-time, Markovian, non-unitary evolution for the system of interest [1, 2].
@@
Besides these theoretical developments, unlike their Hermitian counterparts, the toolbox of computational many-body physics is more limited when it comes to non-unitary dynamics. In this work, we introduce a new method to tackle the quantum dynamics of a non-unitary system. Our approach is based on expanding the evolution operator in Faber polynomials [47, 48]. This numerical approach is a natural generalisation of its Hermitian counterpart, the Chebyshev polynomial method for time evolution [49], which has been the primary choice for efficiently simulating nonequilibrium transport phenomena in both interacting [50] and non-interacting systems [51, 52, 53]. Although time evolution integrators based on Faber polynomials have already been proposed in some works [54, 55, 56], for example, in the simulation of electromagnetic wave propagation through passive media [57, 58], it appears that their full potential remains largely unexplored. Namely, in the simulation of non-Hermitian quantum systems, they could have a significant impact owing to their numerical stability and adjustable accuracy compared to other methods such as integrators based on Runge-Kutta [59] or Trotterization techniques [60].
@@ 2 Non-Unitary Dynamics of Open Quantum Systems
and a series of stochastic quantum jumps at random times, at which the wave function changes discontinuously (see second line of Eq. 1). We note that the non-Hermitian evolution is normalised and state dependent. This is encoded in the last term in the first line of Eq. 1. If one post-selects the quantum trajectories over the records of no click, the dynamics is deterministic and driven by $\mathcal{H}_{\rm nH}$ [33, 32, 31]. Otherwise, if one considers all the trajectories and averages over the measurement outcomes, the conditional density matrix $\rho_{c}(\xi_{t},t)=|\psi(\xi_{t},t)\rangle\langle\psi(\xi_{t},t)|$, i.e.
@@
In both cases the basic building block of the non-unitary dynamics is the evolution driven by a non-Hermitian Hamiltonian. In the next section, we introduce the Faber polynomial method to accurately solve the time evolution governed by a non-unitary Schrödinger equation,
@@ 3 Faber Polynomial Method
where we define |$\ket{\Psi_{n}}=\mathcal{P}_{n}\left(\mathcal{H}\right)\ket{\Psi\left(t_{0}\right)}$. As will be demonstrated, the coefficients $c_{n}(\delta t)$ decrease as the order $n$ increases. Moreover, the states |$\ket{\Psi_{n}}$ are efficiently computed through the recurrence relations that the polynomials satisfy. To compute the subsequent level of the expansion, the main computational task is the application of the system’s Hamiltonian onto a particular state. Consequently, the most demanding operation involves only multiplying the Hamiltonian by a limited group of vectors, leading to a resource usage that increases linearly with $\dim\left(\mathcal{H}\right)$ for sparse matrices or quadratically with $\dim\left(\mathcal{H}\right)$ for dense matrices. Linear scaling is expected for Hamiltonians characterising a system with short-range interactions or hoppings. These principles are exactly those underpinning Kernel Polynomial Methods [64], which has become an essential computational resource in condensed matter physics, particularly for calculating various spectral quantities[65, 66, 67, 68, 69].
@@ 3.1 Warm-Up: Unitary Evolution
When the time evolution is generated by a Hermitian Hamiltonian, one can expand Eq. 5 using Chebyshev polynomials of the first kind (for further details check Tal-Ezer and Kosloff [49]),
@@
where ~$\tilde{\mathcal{H}}=\mathcal{H}/\lambda$ is the rescaled Hamiltonian, $J_{n}(x)$ is the $n^{\text{th}}$ Bessel Function of the first kind, and $T_{n}$ is the $n^{\text{th}}$ Chebyshev polynomial of the first kind. It is necessary to rescale the Hamiltonian so that its eigenvalues fall within the domain of the definition of polynomials, the open interval $\left(-1,1\right)$. It is always possible to do this given that the Hamiltonian is always bounded in finite-size lattice models. Using the recursion relation of the Chebyshev polynomials, the states |$\ket{\Psi_{n}}$ in Eq. 7 are computed on the run using,
@@ 3.2 Non-Unitary Evolution
For a non-Hermitian Hamiltonian, the propagator Eq. 5 is expanded using Faber polynomials  [70, 71] instead. These are a familiar tool in complex analysis, used as a polynomial basis to represent a complex-valued function within the domain $\mathcal{D}$ in which it is analytical. The Faber polynomials are generated by conformal mapping, which maps the complement of a closed disk of radius $\rho$ to the complement of the region containing all the spectra of the Hamiltonian (the domain $\mathcal{D}$). In our work, we assume $\mathcal{D}$ to be an elliptic region containing all the eigenvalues of $\mathcal{H}$. This fits our purposes, as the conformal mapping associated with this shape generates a class of Faber polynomials with a minimum recurrence relation (see the Appendix A for further details). The expansion of Eq. 5 for a non-Hermitian Hamiltonian with this choice reduces to
$\mathcal{U}\left(t\right)=\sum_{n=0}^{+\infty}c_{n}\left(t\right)F_{n}\left(\tilde{\mathcal{H}}\right),\quad c_{n}(t)=e^{-i\lambda t\gamma_{0}}\left(\frac{-i}{\sqrt{\gamma_{1}}}\right)^{n}\;J_{n}\left(2\sqrt{\gamma_{1}}\lambda t\right).$ (10)
@@
[…]
@@
|$\ket{\Psi_{0}}$ =|$=\ket{\Psi\left(t_{0}\right)}$ (12)
|$\ket{\Psi_{1}}$ =(~$=\left(\tilde{\mathcal{H}}-\gamma_{0}\right)\ket{\Psi_{0}}$
|$\ket{\Psi_{2}}$ =(~$=\left(\tilde{\mathcal{H}}-\gamma_{0}\right)\ket{\Psi_{1}}-2\gamma_{1}\ket{\Psi_{0}}$
|$\ket{\Psi_{n+1}}$ =(~$=\left(\tilde{\mathcal{H}}-\gamma_{0}\right)\ket{\Psi_{n}}-\gamma_{1}\ket{\Psi_{n-1}},\quad n>2$
In order to perform one time step of evolution, one simply has to truncate Eq. 10 up to the desired order $N_{p}$, and calculate the associated states |$\ket{\Psi_{n}}$. Through the relations in Eq. 12, one never has to store the Hamiltonian matrix, only needing to store in memory at most the previous two states, |$\ket{\Psi_{n-1}}$ and |$\ket{\Psi_{n}}$, to compute the following term of the expansion |$\ket{\Psi_{n+1}}$. Furthermore, the expansion coefficients can be computed once at the beginning of the algorithm, as they depend only on the chosen time step. Given this, the computation of the state of the system after a time step scales linearly in the number of polynomials and linearly in the Hilbert space dimension (assuming that the Hamiltonian has a sparse representation on the used basis). This scaling can be improved by using parallelisation techniques [68] and making use of the underlying symmetries of the Hamiltonian [72].
@@ 4 Application to Non-Hermitian Gaussian Systems
with Q a unitary $\rm L\times M$ matrix. So the proper-normalised many-body state is obtained by assigning $\text{U}(t+\delta t)=Q$. Even tough, thought in this document we only study particle-number conserving models, this method is easily also applied to non-particle conserving Hamiltonians using similar techniques [75, 20]. Typically, this normalisation step is the most computationally intensive part of the time-evolution process. However, unlike other methods, this step can be executed less frequently since the time-step does not need to be small. Thus, this computationally intensive procedure can be minimised while still ensuring high accuracy in the time integration.
The method of directly evolving the state presents a more cost-effective alternative to approaches that compute the equations of motion for all two-point functions. The latter typically employs conventional ordinary differential equation solvers, such as the fourth-order Runge-Kutta method. Firstly, it necessitates the selection of a short time-step to ensure accurate integration, a requirement not imposed by the Faber algorithm. Secondly, this method involves the evolution of the full correlation matrix, which in practice involves the evolution of $\rm L\cdot\left(L-1\right)/2$ elements. In contrast, the evolution of U requires the evolution of $\rm L\times M$ elements with $\rm M\leq L$. Lastly, the Faber algorithm requires the storage of only two vectors of size $\rm L$ in memory to evolve a given column, while the integration of the correlation matrix demands the storage of four vectors of size $\rm L$ to evolve each of the $\rm L$ columns.
@@ 4.1 Benchmark: Dynamics of the Hatano-Nelson model
where $J$ is a hopping term to the first-neighbour, $\gamma\in\mathbb{R}$ parameterised the left-right imbalance in charge hopping, also called non-reciprocity, and $c^{\dagger}_{\ell}\left(c_{\ell}\right)$ is the creation (annihilation) operator which creates (destroys) a fermion on site $\ell$. Under open boundary conditions (OBC), this non-reciprocal hopping gives rise to a unique phenomenon of non-Hermitian systems, the skin effect [35, 36, 37, 38, 39, 40]. This corresponds to the localization of the single-particle eigenstates at the edges of the system. In addition, the model has a huge sensitivity to the boundary conditions: under periodic boundary conditions (PBC), the single-particle spectrum encircles $E=0$ in the complex plane, with the eigenstates manifesting as delocalised plane waves. In contrast, with OBC, the spectrum is real for $\left|\gamma\right|<J$ and purely imaginary for $\left|\gamma\right|>J$. Additionally, all right single-particle eigenstates show an exponential localisation at the left (right) boundary for positive (negative) $\gamma$ (check Appendix. B for further details).
Figure 2: Left panel: Comparison between the results obtained with the Faber polynomial method and those reported in [34] for the entanglement entropy of half of the chain for a total chain of size $L=100$. Right panel: Entanglement entropy for half of the chain, with $\gamma=-0.05J$ using a different number of polynomials and the time step of $0.1J^{-1}$. We note that we use a symmetric definition for $\gamma$ with regard to [34]
In the following, we benchmark our results using the Faber polynomial method, with those reported by Kawabata et al. [34]. That is, we investigate the dynamics of the entanglement [76] associated with a segment of the chain, denoted $\ell$. This is rigorously derived from the von Neumann entropy of the reduced density matrix($\rho_{\ell}$[77],
@@
$\rho_{\ell}$ is determined by tracing out the complementary degrees of freedom of the subregion $\ell$. However, the Gaussianity of the state allows us to use the standard techniques [78, 79, 80] to perform this computation using the one-particle density matrix restricted to the lattice sites belonging to the region $\ell$, $\mathcal{C}_{n,m\in\ell}=\left\langle c^{\dagger}_{n}c_{m}\right\rangle$. Thus, Eq. 19 for a free fermionic system is simplified to
@@
We see that, consistently with the density plot, a finite (negative) current flows in the bulk of the chain, for sufficiently long systems and long times, while at the boundaries the current vanishes due to the localised charge in the domain walls. This current is a feature of this non-equilibrium steady, given that it is not present in the ground-state of the Hatano-Nelson model. Besides, in contrast to the Hermitian scenario, the density profiles exhibit a spatial gradient within the bulk of the chain. This non-trivial spatial distribution of density arises from the single-particle skin effect. Additionally, in contrast to the Hermitian scenario, non-Hermiticity permits spatial variations in the local charge current while the density profile remains time-independent. This phenomenon is facilitated by the fact that the continuity equation associated to charge conservation in non-Hermitian systems takes the form
@@
Figure 3: Left - Time and spatial dependence of the particle density (Top plot) and charge current (bottom) profile for a system size of 100 sites. Right - Spatial dependence of the particle density (top) and charge current (bottom) in the steady for different system sizes. Other parameters: $\gamma=-0.8J$
Figure 4: Time evolution of the magnetisation profile of the Hatano-Nelson model for different values of the non-Hermitian parameter $\gamma$. The total system length corresponds to $L=256$. The dashed white lines correspond to the effective penetration length.

### v3

@@ 1 Introduction
In recent years, the scientific community has shown a growing interest in elucidating the distinctive characteristics of many-body quantum systems subjected to effective non-unitary dynamics. In quantum mechanics, non-unitary dynamics typically arises when a closed quantum system interacts with an external environment, leading to dissipation, decoherence, or wave function collapse that disrupts the usual unitary Schrödinger evolution. While a fully microscopic description of both environment and system is a daunting task, in many cases of experimental and theoretical interest, one can assume the dynamics of the environment to be sufficiently fast, allowing for the derivation of a local-in-time, Markovian, non-unitary evolution for the system of interest [1, 2].
@@
Besides these theoretical developments, unlike their Hermitian counterparts, the toolbox of computational many-body physics is more limited when it comes to non-unitary dynamics. In this work, we introduce a new method to tackle the quantum dynamics of a non-unitary system. Our approach is based on expanding the evolution operator in Faber polynomials [47, 48]. This numerical approach is a natural generalisation of its Hermitian counterpart, the Chebyshev polynomial method for time evolution [49], which has been the primary choice for efficiently simulating nonequilibrium transport phenomena in both interacting [50] and non-interacting systems [51, 52, 53]. Although time evolution integrators based on Faber polynomials have already been proposed in some works [54, 55, 56], for example, in the simulation of electromagnetic wave propagation through passive media [57, 58], it appears that their full potential remains largely unexplored. Namely, in the simulation of non-Hermitian quantum systems, they could have a significant impact owing to their numerical stability and adjustable accuracy compared to other methods such as integrators based on Runge-Kutta [59] or Trotterization techniques [60]. Furthermore, the use of Faber Polynomials could replace current techniques [61, 62, 63] that still employ Chebyshev polynomials but need to rely on hermitization. This results in the need to operate within a vector space that has twice the original dimensionality.
@@ 2 Non-Unitary Dynamics of Open Quantum Systems
and a series of stochastic quantum jumps at random times, at which the wave function changes discontinuously (see second line of Eq. (1)). We note that the non-Hermitian evolution is normalised and state dependent. This is encoded in the last term in the first line of Eq. (1). If one post-selects the quantum trajectories over the records of no click, the dynamics is deterministic and driven by $\mathcal{H}_{\rm nH}$ [33, 32, 31]. Otherwise, if one considers all the trajectories and averages over the measurement outcomes, the conditional density matrix $\rho_{c}(\xi_{t},t)=|\psi(\xi_{t},t)\rangle\langle\psi(\xi_{t},t)|$, i.e.
@@
In both cases, the basic building block of the non-unitary dynamics is the evolution driven by a non-Hermitian Hamiltonian. In the next section, we introduce the Faber polynomial method to accurately solve the time evolution governed by a non-unitary Schrödinger equation.
@@ 3 Faber Polynomial Method
Hamiltonianwhere we define |$\ket{\Psi_{n}}=\mathcal{P}_{n}\left(\mathcal{H}\right)\ket{\Psi\left(t_{0}\right)}$. As will be demonstrated, the coefficients $c_{n}(\delta t)$ decrease as the order $n$ increases. Moreover, the states |$\ket{\Psi_{n}}$ are efficiently computed through the recurrence relations that the polynomials satisfy. To compute the subsequent level of the expansion, the main computational task is the application of the system’s Hamiltonian onto a particular state. Consequently, the most demanding operation involves only multiplying the Hamiltonian by a limited group of vectors, leading to a resource usage that increases linearly with $\dim\left(\mathcal{H}\right)$ for sparse matrices or quadratically with $\dim\left(\mathcal{H}\right)$ for dense matrices. Linear scaling is expected for Hamiltonians characterising a system with short-range interactions or hoppings. These principles are exactly those underpinning Kernel Polynomial Methods [67], which has become an essential computational resource in condensed matter physics, particularly for calculating various spectral quantities[68, 69, 70, 71, 72]. There has also been established a Non-Hermitian Kernel Polynomial Method [61, 62, 63], designed for computing dynamic correlators or spectral functions of the Liouvillian superoperator (or of an effective non-Hermitian Hamiltonian), which still uses Chebyshev polynomials through hermitization techniques.
@@ 3.1 Warm-Up: Unitary Evolution
When the time evolution is generated by a Hermitian Hamiltonian, one can expand Eq. (5) using Chebyshev polynomials of the first kind (for further details check Tal-Ezer and Kosloff [49]),
@@
where ~$\tilde{\mathcal{H}}=\mathcal{H}/\lambda$ is the rescaled Hamiltonian, $J_{n}(x)$ is the $n^{\text{th}}$ Bessel Function of the first kind, and $T_{n}$ is the $n^{\text{th}}$ Chebyshev polynomial of the first kind. It is necessary to rescale the Hamiltonian so that its eigenvalues fall within the domain of the definition of polynomials, the open interval $\left(-1,1\right)$. It is always possible to do this given that the Hamiltonian is always bounded in finite-size lattice models. Using the recursion relation of the Chebyshev polynomials, the states |$\ket{\Psi_{n}}$ in Eq. (7) are computed on the run using,
@@ 3.2 Non-Unitary Evolution
For a non-Hermitian Hamiltonian, the propagator Eq. (5) is expanded using Faber polynomials  [73, 74] instead. These are a familiar tool in complex analysis, used as a polynomial basis to represent a complex-valued function within the domain $\mathcal{D}$ in which it is analytical. The Faber polynomials are generated by conformal mapping, which maps the complement of a closed disk of radius $\rho$ to the complement of the region containing all the spectra of the Hamiltonian (the domain $\mathcal{D}$). In our work, we assume $\mathcal{D}$ to be an elliptic region containing all the eigenvalues of $\mathcal{H}$. This fits our purposes, as the conformal mapping associated with this shape generates a class of Faber polynomials with a minimum recurrence relation (see the Appendix A for further details). The expansion of Eq. (5) for a non-Hermitian Hamiltonian with this choice reduces to
$\mathcal{U}\left(t\right)=\sum_{n=0}^{+\infty}c_{n}\left(t\right)F_{n}\left(\tilde{\mathcal{H}}\right),\quad c_{n}(t)=e^{-i\lambda t\gamma_{0}}\left(\frac{-i}{\sqrt{\gamma_{1}}}\right)^{n}\;J_{n}\left(2\sqrt{\gamma_{1}}\lambda t\right),$ (10)
@@
$\lambda=\frac{\left(\ell^{2/3}+p^{2/3}\right)^{3/2}}{2},\quad\gamma_{1}=\frac{\left(\tilde{p}^{2/3}+\tilde{\ell}^{2/3}\right)\left(\tilde{p}^{4/3}-\tilde{\ell}^{4/3}\right)}{4\lambda},$ (11)
@@
|$\ket{\Psi_{0}}$ =|$=\ket{\Psi\left(t_{0}\right)},$ (12)
|$\ket{\Psi_{1}}$ =(~$=\left(\tilde{\mathcal{H}}-\gamma_{0}\right)\ket{\Psi_{0}},$
|$\ket{\Psi_{2}}$ =(~$=\left(\tilde{\mathcal{H}}-\gamma_{0}\right)\ket{\Psi_{1}}-2\gamma_{1}\ket{\Psi_{0}},$
|$\ket{\Psi_{n+1}}$ =(~$=\left(\tilde{\mathcal{H}}-\gamma_{0}\right)\ket{\Psi_{n}}-\gamma_{1}\ket{\Psi_{n-1}},\quad n>2.$
In order to perform one time step of evolution, one simply has to truncate Eq. (10) up to the desired order $N_{p}$, and calculate the associated states |$\ket{\Psi_{n}}$. Through the relations in Eq. (12), one never has to store the Hamiltonian matrix, only needing to store in memory at most the previous two states, |$\ket{\Psi_{n-1}}$ and |$\ket{\Psi_{n}}$, to compute the following term of the expansion |$\ket{\Psi_{n+1}}$. Furthermore, the expansion coefficients can be computed once at the beginning of the algorithm, as they depend only on the chosen time step. Given this, the computation of the state of the system after a time step scales linearly in the number of polynomials and linearly in the Hilbert space dimension (assuming that the Hamiltonian has a sparse representation on the used basis). This scaling can be improved by using parallelisation techniques [71] and making use of the underlying symmetries of the Hamiltonian [75].
@@ 4 Application to Non-Hermitian Gaussian Systems
with Q a unitary $\rm L\times M$ matrix. So the proper-normalised many-body state is obtained by assigning $\text{U}(t+\delta t)=Q$. Although this document focuses on particle-number conserving models, the method can also be easily applied to non-particle conserving Hamiltonians using similar techniques [78, 20]. Typically, this normalisation step is the most computationally intensive part of the time-evolution process. However, unlike other methods, this step can be executed less frequently since the time-step does not need to be small. Thus, this computationally intensive procedure can be minimised while still ensuring high accuracy in the time integration.
The method of directly evolving the state presents a more cost-effective alternative to approaches that solve the equations of motion for all two-point functions. The latter typically employs conventional ordinary differential equation solvers, such as the fourth-order Runge-Kutta method. Firstly, Runge-Kutta methods are only accurate up to a given order. As such, to ensure an accurate time integration, it is necessary to select a short time-step in comparison to the energy scales of the problem, a requirement not imposed by the Faber algorithm. Secondly, the Runge-Kutta method involves the evolution of the full correlation matrix, which in practice involves the evolution of $\rm L\cdot\left(L-1\right)/2$ elements. In contrast, the evolution of U requires the evolution of $\rm L\times M$ elements with $\rm M\leq L$. Finally, the Faber algorithm only needs to store two vectors of length $\rm L$ in memory to evolve a given column, while the integration of the correlation matrix demands the storage of four vectors of size $\rm L$ to evolve each of the $\rm L$ columns.
The Faber algorithm is also more beneficial compared to methods that use the exponential of the Hamiltonian. Firstly, one would need to store the exponentiated matrix for repeated application without recalculations. In our approach, it is only necessary to store the system state, since the Hamiltonian matrix is never kept in memory. Additionally, the exponentiated matrix is typically a dense matrix, leading to a computational cost that scales quadratically with its size when multiplied by a vector. Conversely, the Faber Polynomial technique takes advantage of the possible sparcities of the Hamiltonian matrix, making the computational cost linear with respect to the dimension of the single-particle Hilbert space.
@@ 4.1 Benchmark: Dynamics of the Hatano-Nelson model
where $J$ is a hopping term to the first-neighbour, $\gamma\in\mathbb{R}$ parameterised the left-right imbalance in charge hopping, also called non-reciprocity, and $c^{\dagger}_{\ell}\left(c_{\ell}\right)$ is the creation (annihilation) operator which creates (destroys) a fermion on site $\ell$. Under open boundary conditions (OBC), this non-reciprocal hopping gives rise to a unique phenomenon of non-Hermitian systems, the skin effect [35, 36, 37, 38, 39, 40]. This corresponds to the localisation of the single-particle eigenstates at the edges of the system. In addition, the model has a huge sensitivity to the boundary conditions: under periodic boundary conditions (PBC), the single-particle spectrum encircles $E=0$ in the complex plane, with the eigenstates manifesting as delocalised plane waves. In contrast, with OBC, the spectrum is real for $\left|\gamma\right|<J$ and purely imaginary for $\left|\gamma\right|>J$. Additionally, all right single-particle eigenstates show an exponential localisation at the left (right) boundary for positive (negative) $\gamma$ (check Appendix B for further details).
Figure 2: Left panel: Comparison between the results obtained with the Faber polynomial method and those reported in [34] for the entanglement entropy of half of the chain for a total chain of size $L=100$. Right panel: Entanglement entropy for half of the chain, with $\gamma=-0.05J$ using a different number of polynomials and the time step of $0.1J^{-1}$. We note that we use a symmetric definition for $\gamma$ with regard to [34].
In the following, we benchmark our results using the Faber polynomial method, with those reported by Kawabata et al [34]. That is, we investigate the dynamics of the entanglement [79] associated with a segment of the chain, denoted $\ell$. This is rigorously derived from the von Neumann entropy of the reduced density matrix($\rho_{\ell}$[80],
@@
$\rho_{\ell}$ is determined by tracing out the complementary degrees of freedom of the subregion $\ell$. However, the Gaussianity of the state allows us to use the standard techniques [81, 82, 83] to perform this computation using the one-particle density matrix restricted to the lattice sites belonging to the region $\ell$, $\mathcal{C}_{n,m\in\ell}=\left\langle c^{\dagger}_{n}c_{m}\right\rangle$. Thus, Eq. (19) for a free fermionic system is simplified to
@@
We see that, consistently with the density plot, a finite (negative) current flows in the bulk of the chain, for sufficiently long systems and long times, while at the boundaries the current vanishes due to the localised charge in the domain walls. This current is a feature of this non-equilibrium steady-state, given that it is not present in the ground-state of the Hatano-Nelson model. Besides, in contrast to the Hermitian scenario, the density profiles exhibit a spatial gradient within the bulk of the chain. This non-trivial spatial distribution of density arises from the single-particle skin effect. Additionally, in contrast to the Hermitian scenario, non-Hermiticity permits spatial variations in the local charge current while the density profile remains time-independent. This phenomenon is facilitated by the fact that the continuity equation associated to charge conservation in non-Hermitian systems takes the form
@@
Figure 3: Left - Time and spatial dependence of the particle density (top plot) and charge current (bottom) profile for a system size of 100 sites. Right - Spatial dependence of the particle density (top) and charge current (bottom) in the steady for different system sizes. Other parameters: $\gamma=-0.8J$.
Figure 4: Time evolution of the particle density profile of the Hatano-Nelson model for different values of the non-Hermitian parameter $\gamma$. The total system length corresponds to $L=256$. The dashed white lines correspond to the effective penetration length.